# Obstruction to Broken Symmetries in Topological Flat Bands

### v1

@@ Abstract
One well-known example of such is the BCS-type superconductivity (SC) in a magnetic Weyl semimetal, where pairing order parameter between two Fermi surfaces (FSs) enclosing two Weyl points of opposite chirality must be nodal [15, 16]. Interestingly, there is a topological obstruction to a gapped/nodeless SC state due to the Chern number of paired electrons on the FSs. Inspired by this result, topological obstruction to gapped superconductors have been discussed for pairings between electrons in normal states with nontrivial $\mathbb{Z}_{2}$ topology [17], fragile topology [18, 19], and delicate topology [20]. Moreover, topologically obstructed gapless/nodal charge order has also been studied for electrons in topological bands with nonzero Chern numbers [21, 22]. The fundamental idea behind these works is to establish a connection between the vorticity of the order parameters and the topological invariants of electronic bands. As a result, the presence of nontrivial topological invariants indicates the existence of nonvanishing vorticities of the order parameters in momentum space, and thus zero points of the order parameters. These arguments rely on the mean-field picture by analyzing the geometry and topology of non-interacting Bloch wavefunctions, which breaks down in the presence of strong interactions. Therefore, a natural question is: do the topological obstruction to gapped SC and charge orders persist in strongly correlated systems? Besides bearing the theoretical importance, addressing this question is also helpful for understanding ordered phases in realistic materials, given that the topologically obstructed SC order and inter-valley coherent order have been proposed to occur in flat bands of Moire systems based on band theory [19, 23, 22], where the interactions are strong compared to the small band width.
@@
Discussions— In this work, we have established an LSM-type topological obstruction to trivially gapped SC and charge orders in strongly correlated flat Chern bands at half filling. Our results reveal an intriguing connection between long-range orders and topological orders, i.e., the presence of certain symmetric, gapped long-range orders is necessarily accompanied by topological orders. Since the Landau-type long-range orders are much easier to detect in experiments than topological orders, the topological obstructions discussed in this work may serve as a indicator for the presence of topological orders. Importantly, the flat Chern bands discussed in this work can be realized either in 2D Morie systems [37, 38, 22, 39] and/or on the Fermi surface of 3D topological semimetals such as Weyl semimetals [16, 40, 41], for which the $U_{v}(1)$ symmetry are indeed realized as a valley-$U(1)$ symmetry. Both topological semimetals and 2D Morie systems are important material platforms in experiments, and thus our results can provide new insights to the study of SC and/or inter-valley coherent orders in these systems.
@@ Acknowledgements.
Acknowledgments: We thank Jiabin Yu, Ya-Hui Zhang, and Rui-Xing Zhang for insightful discussions. This research was primarily supported by the Center for Emergent Materials, an NSF MRSEC, under award number DMR-2011876.

### v2

@@ Abstract
One well-known example of such is the BCS-type superconductivity (SC) in a magnetic Weyl semimetal, where pairing order parameter between two Fermi surfaces (FSs) enclosing two Weyl points of opposite chirality must be nodal [15, 16]. Interestingly, there is a topological obstruction to a gapped/nodeless SC state due to the Chern number of paired electrons on the FSs. Inspired by this result, topological obstruction to gapped superconductors have been discussed for pairings between electrons in normal states with nontrivial $\mathbb{Z}_{2}$ topology [17], fragile topology [18, 19], and delicate topology [20]. Moreover, topologically obstructed gapless/nodal charge order has also been studied for electrons in topological bands with nonzero Chern numbers [21, 22]. The fundamental idea behind these works is to establish a connection between the vorticity of the order parameters and the topological invariants of electronic bands. As a result, the presence of nontrivial topological invariants indicates the existence of nonvanishing vorticities of the order parameters in momentum space, and thus zero points of the order parameters. These arguments rely on the mean-field picture by analyzing the geometry and topology of non-interacting Bloch wavefunctions, which breaks down in the presence of strong interactions. Therefore, a natural question is: do the topological obstruction to gapped SC and charge orders persist in strongly correlated systems? Besides bearing the theoretical importance, addressing this question is also helpful for understanding ordered phases in realistic materials, given that the topologically obstructed SC order and inter-valley coherent (IVC) order have been proposed to occur in flat bands of Moire systems based on band theory [19, 23, 22], where the interactions are strong compared to the small band width.
@@
Discussions— In this work, we have established an LSM-type topological obstruction to trivially gapped SC and charge orders in strongly correlated flat Chern bands at half filling. Our results reveal an intriguing connection between long-range orders and topological orders, i.e., the presence of certain symmetric, gapped long-range orders is necessarily accompanied by topological orders. Since the Landau-type long-range orders are much easier to detect in experiments than topological orders, the topological obstructions discussed in this work may serve as a indicator for the presence of topological orders. Importantly, the flat Chern bands discussed in this work can be realized either in 2D Morie systems [37, 38, 22, 39] and/or on the Fermi surface of 3D topological semimetals such as Weyl semimetals [16, 40, 41], for which the $U_{v}(1)$ symmetry are indeed realized as a valley-$U(1)$ symmetry. We emphasize that the emergent chiral symmetry, arising from projecting the Hilbert space onto flat bands, is vital for the topological obstruction. If this symmetry is broken, a gapped incommensurate valley coherence (IVC) order can emerge. In twisted bilayer-graphenes, the finite width and nontrivial dispersion of the flat Chern bands will generally break the emergent chiral symmetry. As a result, incommensurate Kekulé orders have been theoretically proposed [42, 43, 44] experimentally confirmed [45]. The emergent chiral symmetry of flat bands in real materials should be considered as an approximate symmetry. The flatter the bandwidth relative to the interaction strength, the more accurate this approximation becomes. Therefore, we expect the topological obstruction discussed in our paper would be more influential in systems with flatter Chern bands, e.g. in twisted MoTe2 homobilayers [46, 47, 48, 49], where the breaking of the emergent chiral symmetry is expected to be weaker and thus it is harder to have a trivially gapped IVC order.
@@ Acknowledgements.
Acknowledgments: We thank Jiabin Yu, Ya-Hui Zhang, and Rui-Xing Zhang for their insightful discussions. We also extend our gratitude to Siddharth A. Parameswaran and Nemin Wei for their valuable discussions on materials. This research was primarily supported by the Center for Emergent Materials, an NSF MRSEC, under award number DMR-2011876.